OpenStudy (sebaxtiangb):

who knows trigonometry? i need help proving cot(a+b)={ cot(a)cot(b) - 1}/cot(b)cot(a)

OpenStudy (jango_in_dtown):

@sebaxtiangb

OpenStudy (jango_in_dtown):

start from rhs. and put cot a and cot b in the from cos a/sin a and cos b/ sin b.

OpenStudy (jango_in_dtown):

then after a stage you will get (cos a cos b-sin a sin b)/(sin a cos b +cos a sin b)=cos (a+b)/sin(a+b)=cot (a+b)

OpenStudy (sebaxtiangb):

alright thanks